### v1

# VLMQ: Efficient Post-Training Quantization for Large Vision-Language Models via Hessian Augmentation

Abstract
[…]
1 Introduction
Large language models (LLMs) (Bubeck et al. 2023; Touvron et al. 2023a, b; DeepSeek-AI 2025; Jiang et al. 2023) have demonstrated exceptional advancements across diverse natural language processing tasks, leading to an increased emphasis on developing vision-language models (VLMs) (Wang et al. 2024b; Bai et al. 2025; Li et al. 2024b; Chen et al. 2024d) that process multi-modal inputs, including texts, images, and videos. Despite their impressive capabilities, the unprecedented scaling in the model size complicates their deployment in practical resource-limited contexts. Quantization methods have provided an effective solution by converting full-precision weights and activations (e.g., FP16/BF16) into reduced-precision formats (e.g., INT8/INT4), thereby markedly lowering memory footprint and computation complexity at the expense of minor performance degradation.
Figure 1: Baseline vs. VLMQ. The diagonal matrix in VLMQ represents the importance factors. The reported accuracy is from 2-bit Qwen2-VL-7B-Instruct (Wang et al. 2024b) quantized by GPTQ and VLMQ.
Post-training quantization (PTQ) has emerged as a prevalent approach for deploying large-scale models, owing to its minimal computational overhead and the ability to bypass the costly fine-tuning or retraining process. Substantial research efforts have been directed toward designing advanced PTQ methods tailored LLMs. Such approaches typically aim to refine the distributions of weights and activations through various strategies, including equivalent transformations (Lin et al. 2023; Xiao et al. 2023), Hessian-based error compensation (Frantar and Alistarh 2022; Li et al. 2025b), and incoherence processing (Ashkboos et al. 2024; Hu et al. 2025). While significant progress has been achieved in applying PTQ to LLMs, their extension to VLMs has not yet been adequately investigated. Recent pioneering research efforts, such as MBQ (Li et al. 2025a), MQuant (Yu et al. 2025), and QSLAW (Xie et al. 2024), emphasize the modality imbalance challenge and propose dedicated importance-aware strategies to enhance the performance of quantized VLMs. However, these methods either require expensive parameter fine-tuning (Xie et al. 2024), specialized manipulation at inference time (Yu et al. 2025), or rely on a suboptimal grid search (Li et al. 2025a), failing to offer an efficient and effective solution for both calibration and inference.
Figure 2: PCA-based Hessian feature analysis with activations (4096 points) extracted from the pre-attention trace of the $20$-th transformer layer in Qwen2-VL-7B-Instruct. The left two subfigures depict the Hessian feature distributions constructed from text-only and mixed text-vision activations, respectively. The right three subfigures visualize the distributions with varying token-level importance factors. Light red/green and dark red/green points denote tokens classified as important or unimportant ones. Reported average accuracy is under INT3 quantization across eight vision-language benchmarks.
In this work, we identify a modality disproportion (i.e., limited text tokens vs. excessive and redundant vision tokens) in VLMs (Chen et al. 2024c; Xing et al. 2024; Alvar et al. 2025). We refer to redundancy as vision over-representation. Existing Hessian-based PTQ methods, such as GPTQ (Frantar et al. 2022), optimize the layer-wise reconstruction loss by treating all tokens equally. In these methods, the Hessian matrix that serves as a crucial guide for weight quantization is estimated by uniformly aggregating token contributions, regardless of their informativeness or importance. This token-agnostic formulation inherently biases the Hessian toward dominant yet redundant vision features, leading to sub-optimal weight updates. As a result, directly applying these methods to VLMs results in severe performance degradation.
To mitigate the side effects of vision over-representation, we propose VLMQ, an importance-aware calibration framework that optimizes a refined objective. As illustrated in Figure 1, VLMQ modifies the standard Hessian into an enhanced variant by assigning higher importance to salient tokens and lower importance to redundant ones. It yields a closed-form weight update while capturing token-level importance via the augmented Hessian. VLMQ is fully compatible with existing Hessian-based PTQ frameworks and readily inherits efficient tricks such as Cholesky decomposition. To derive effective importance factors, we first establish a theoretical connection between block-level loss perturbation and token-level quantization error. Based on this insight, we propose to perform a single lightweight block-wise backpropagation approach to derive layer-wise gradient-driven importance factors efficiently.
The core contributions are summarized as follows:
  • We identify a critical mismatch between the vision redundancy in VLMs and the token-agnostic objective of existing Hessian-based PTQ methods, which treat all tokens equally during quantization. We posit that this overlooked vision redundancy distorts the Hessian estimation, thereby explaining the poor performance of prior LLM PTQ methods when applied to VLMs.
  • We propose a PTQ scheme that leverages an enhanced Hessian derived from an importance-aware objective, which explicitly accounts for the redundancy among vision tokens and supports high degrees of parallelism.
  • We theoretically connect block-wise loss perturbation and first-order token-level activation errors, which motivates the use of a single lightweight block-wise backward pass to derive gradient-driven importance factors. This design achieves both efficiency and effectiveness in capturing the importance of tokens.
  • Through extensive experimentation, we demonstrate the superior performance of our approach for VLMs, especially under ultra-low-bit quantization settings. For example, our method achieves state-of-the-art (SOTA) results and achieves up to 16.45% accuracy improvement.
2 Preliminary
Notation. We follow most notations in GPTAQ (Li et al. 2025b). Specifically, general model parameters including $N$, $C_{i}$, $C_{o}$, and $\mathcal{V}$} denote the sequence length, in-channel number (hidden size), out-channel number (intermediate size), and the vocabulary set. We adopt the bold lowercase and uppercase letters to represent row-vectors and matrices, respectively. The letters with a hat symbol (e.g., $\hat{{\mathbf{x}}}$ or $\hat{{\mathbf{W}}}$) represent the quantized weights or activations noised by quantization errors. The linear operation is described as ${\mathbf{Y}}={\mathbf{W}}{\mathbf{X}}$ where ${\mathbf{W}}\in{\mathbb{R}}^{C_{o}\times C_{i}}$ and ${\mathbf{X}}\in{\mathbb{R}}^{C_{i}\times N}$ represent the weight and activation. The indexing rule is that ${\mathbf{W}}_{j,:}$ and ${\mathbf{W}}_{:,j}$ indicate the $j$-th row and $i$-th column of matrix $\mathbf{W}$} respectively. A negative index signifies the removal of a row in a matrix (e.g., ${\mathbf{X}}_{-1}\in{\mathbb{R}}^{(C_{i}-1)\times N}$).
VLM architecture. The advanced development of VLMs (Wang et al. 2024b; Bai et al. 2025; Chen et al. 2024d; Li et al. 2024b) handles data in different modalities, including text, image, and video. The VLMs comprise three key components: a visual encoder, a vision-text projector, and a Language Model (LM) backbone. The mainstream decoder-only LM backbones output a distribution ${\mathbf{P}}=\left(p_{1},\;p_{2},\;...,\;p_{N-1}\right)\in{\mathbb{R}}^{N\times|{\mathcal{V}}|}$ and refer to it to generate predictions via diverse decoding strategies (Stern, Shazeer, and Uszkoreit 2018; Xia et al. 2022; Kim et al. 2022).
We define input for each decoding layer in the above-mentioned LM backbone as activation ${\mathbf{X}}\in{\mathbb{R}}^{C_{i}\times N}$, where $C_{i}$ and $N$ are the model dimension (i.e., the input channel number) and sequence length, respectively. The activation transforms across $L$ decoding layers in LM backbones. The attention stream and feed-forward network stream within one decoding layer are defined as
$\texttt{Attn}({\mathbf{X}})$ $={\mathbf{X}}+\texttt{MHSA}({\mathbf{X}}),$ (1)
$\texttt{MLP}({\mathbf{X}})$ $={\mathbf{X}}+\texttt{FFN}({\mathbf{X}}),$ (2)
where $\texttt{MHSA}(\cdot)$ describes the multi-head self-attention operation with Q/K/V/O projections and $\texttt{FFN}(\cdot)$ represents the feed forward data flow with Up/Gate/Down projections. We omit the layer normalization involved in decoding layers for simplicity.
3 Related Work
PTQ for VLMs. The PTQ becomes challenging for VLMs due to dissimilar features and distributions between text and vision modality, and remains underexplored. MQuant (Yu et al. 2025) proposes modality-specific static quantization and attention-invariant flexible switching to address the modality gap and visual outliers. MBQ (Li et al. 2025a) observes that the error-sensitivity varies across text and vision tokens, and a corresponding biased objective is formulated to determine the optimal channel-wise equalization scales. QSLAW (Xie et al. 2024) determines the quantization step size via a learning-based method and proposes importance-aware warmup to alleviate the overfitting issue. Unlike these methods concentrating on modality difference, Q-VLM (Wang et al. 2024a) captures cross-layer dependency in VLMs and partitions blocks based on entropy.
OBQ, GPTQ, and GPTAQ. Hessian information is widely used as a calibration guide to compensate for quantization errors. OBQ calibrates the quantization of linear layers in models by iteratively 1) quantizing entries that introduce minimal loss, and 2) updating the remaining unquantized ones with the guidance of Hessian ${\mathbf{H}}={\mathbf{X}}{\mathbf{X}}^{\top}$. GPTQ (Frantar et al. 2022) extends this paradigm by formulating fixed-order entry quantization in a parallel manner, lazy block updating and Cholesky reformulation to achieve efficient calibration. GPTAQ (Li et al. 2025b) further extends GPTQ by optimizing an asymmetric objective:
[…]
where $^{\mathbf{r}}={\mathbf{w}}{\mathbf{X}}-{\mathbf{w}}\hat{{\mathbf{X}}}$ denotes the activation residual and $\Delta{\mathbf{w}}=\hat{{\mathbf{w}}}-{\mathbf{w}}$ is the weight perturbation. The optimal solution is derived via Lagrangian formulation:
$\Delta{\mathbf{w}}=\frac{\left(\hat{{\mathbf{w}}}_{q}-{\mathbf{w}}_{q}\right)}{{\mathbf{H}}_{qq}^{-1}}\cdot{\mathbf{H}}_{q:,}^{-1}+{\mathbf{r}}{\mathbf{X}}^{\top}{\mathbf{H}}_{-q}^{-1},$ (4)
where $q$ is the index of the quantized weight element. GPTAQ thus enables accurate and efficient quantization with better preservation of model functionality.
4 Motivation
Figure 3: Pipeline of computing importance factors. The “Forward” module illustrates the quantization dataflow across decoding layers, where a breakpoint is set at the output of each attention module to compute the local loss ${\mathcal{L}}_{\text{Block}}$ and trigger a localized backward pass. The “Backward” module details the internal operations within an attention block, where gradients of each linear projection output are cached to derive token-level importance factors. The “Process Grad” module summarizes how we formulate importance factors from raw gradients.
Qwen2-VL-7B-Instruct-INT3
Calib Modality LI Ratio DocVQA Acc
Text-only - 86.86%
Text-vision 0% 88.09%
Text-vision 25% 88.19%
Text-vision 50% 88.48%
Text-vision 75% 87.87%
Text-vision (Ours) Fine-grained 88.90%
Table 1: Statistics under controlled settings where a random subset of vision tokens are manually assigned as low-importance (LI) tokens. denotes the percentage of vision tokens whose importance values are set to 0.01, while the remaining tokens retain an importance value of 1.0.
Visual Over-Representation
The visual over-representation problem is a widely recognized issue in VLMs (Chen et al. 2024c; Xing et al. 2024; Alvar et al. 2025), and recent studies have explored vision token reduction to accelerate inference while preserving performance. However, existing Hessian-based PTQ methods, such as GPTQ (Frantar et al. 2022), fail to exploit this redundancy. They construct the Hessian by uniformly aggregating all token contributions that are derived from a standard optimization objective. As a result, these token-agnostic methods often struggle when applied to VLMs, where redundant vision tokens dominate the activation space. This limitation motivates the need for importance-aware quantization strategies that explicitly differentiate informative vision tokens from redundant ones.
Pilot Study
Pilot study on vision-role in quantization. To further investigate the role played by vision tokens in quantization performance, we conduct a pilot study that reveals how vision tokens influence Hessian estimation and model accuracy. We observe that 1) it is essential to incorporate vision tokens for effective quantization in VLMs, yet 2) an excessive number of redundant vision tokens may lead to degraded performance. Assigning redundant tokens with lower importance can mitigate the resulting performance degradation. To validate this insight, we benchmark INT3-quantized LLMs on DocVQA (Mathew, Karatzas, and Jawahar 2021) by randomly down-weighting a portion of vision tokens using a low importance factor. As shown in Table 1, model performance peaks when 50% of vision tokens are marked as low-importance (LI), which indicates that a balanced level of visual input is critical for optimal quantization.
Explanation of the pilot study. We attribute the observed performance fluctuation to the Hessian feature shift, as visualized in Figure 2. The leftmost subfigure shows that text-only tokens yield a compact and centralized distribution in the dominant component space. Introducing vision tokens, as in the second subfigure, diversifies the calibration distribution and helps mitigate the quantization–inference gap. However, due to the visual over-representation, excessive inclusion introduces the risk of Hessian bias toward redundant visual features.
We further color-code token points by their importance factors $\mathbf{G}$} (defined later) in the right three subfigures. The evident imbalance in importance density across modalities risks skewing the Hessian toward redundant visual margins, rather than concentrating on the more informative central regions. These findings highlight the necessity of fine-grained importance-aware quantization to counteract redundancy-induced bias.
Overall, our pilot study highlights the necessity of incorporating visual input, whereas overabundance impairs quantization quality. The performance peak observed with a moderate down-weighting of vision tokens validates the importance of balanced visual input. These insights motivate the need for a fine-grained, importance-aware quantization strategy that selectively emphasizes salient vision tokens while mitigating the side effects of redundant ones.
5 VLMQ
To address the importance discrepancy across tokens, we propose an accurate PTQ method for VLMs by optimizing an importance-aware objective, dubbed VLMQ. Specifically, 1) to capture the varying salience across tokens, we refine the quantization objective to an importance-aware one, yielding an enhanced Hessian. Then, 2) we establish a theoretical link between block-wise loss perturbation and layer-wise output errors, enabling efficient computation of gradient-driven importance scores via a single block-wise backward pass.
Importance-Aware Objective
Although Equation 4 yields the optimal quantized weight $\hat{{\mathbf{w}}}$ by minimizing the squared error, it inherently assumes equal contribution across all tokens in ${\mathbf{Z}}={\mathbf{W}}{\mathbf{X}}$. Such equal treatment may introduce bias due to the vision over-representation. To address this limitation, we introduce a token-level importance weighting matrix ${\mathbf{G}}\in{\mathbb{R}}^{N\times N}$, where $\mathbf{G}$} is diagonal with the $i$-th diagonal element representing the importance assigned to output token ${\mathbf{Z}}_{:,i}$. Incorporating importance factors into the objective leads to a refined formulation:
[…]
$\quad\text{s.t.}\ \Delta{\mathbf{w}}{\mathbf{e}}_{q}^{\top}+{\mathbf{w}}_{q}-\hat{{\mathbf{w}}}_{q}=0$
where ${\mathbf{e}}_{q}$ is the one-hot vector with $q$-th element being 1. We formulate its Lagrangian as:
[…]
which gives the corresponding optimal weight updating as
$\Delta{\mathbf{w}}$ $=\frac{\left(\hat{{\mathbf{w}}}_{q}-{\mathbf{w}}_{q}\right)}{\left[{\mathbf{X}}{\mathbf{G}}{\mathbf{X}}^{\top}\right]_{qq}^{-1}}\cdot\widetilde{{\mathbf{H}}}_{q:,}^{-1}+{\mathbf{r}}{\mathbf{G}}{\mathbf{X}}^{\top}\left[{\mathbf{X}}{\mathbf{G}}{\mathbf{X}}^{\top}\right]_{-q}^{-1}$ (7)
$=\frac{\left(\hat{{\mathbf{w}}}_{q}-{\mathbf{w}}_{q}\right)}{\widetilde{{\mathbf{H}}}_{qq}^{-1}}\cdot\widetilde{{\mathbf{H}}}_{q:,}^{-1}+\tilde{{\mathbf{r}}}\widetilde{{\mathbf{X}}}^{\top}\widetilde{{\mathbf{H}}}_{-q}^{-1},$
where $\widetilde{{\mathbf{H}}}={\mathbf{X}}{\mathbf{G}}{\mathbf{X}}^{\top}$, $\tilde{{\mathbf{r}}}={\mathbf{r}}\sqrt{\mathbf{G}}$ and $\widetilde{{\mathbf{X}}}={\mathbf{X}}\sqrt{\mathbf{G}}$ indicate the importance-aware Hessian, residual, and activation matrix yielded from the refined objective (Equation 5). Our importance-aware weight update presented in Equation 7 remains formally consistent with the original GPTAQ (Li et al. 2025b) formulation. Distinct from BoA (Kim et al. 2024), which attributes importance at the channel level, our approach assigns importance scores at the token-level activation granularity. This design is orthogonal to the parallelization strategies proposed in GPTQ (Frantar et al. 2022) and GPTAQ (Li et al. 2025b). The compatibility allows us to directly inherit efficiency tricks (e.g., Cholesky reformulation) proposed therein.
Gradient-Driven Importance Factors
Figure 4: Visualization of importance factors of k_proj and v_proj outputs.
Figure 3 illustrates the overall process of computing gradient-driven importance factors, which involves three key stages: Forward, Backward, and Gradient Processing. We begin by introducing the forward pass and defining the block-wise target. Next, we theoretically justify the use of gradients for importance estimation and describe how gradient caching is performed during backpropagation. Finally, we present the formal procedure for computing token-level importance factors from the raw gradients.
Forward. During Forward, we place activation hooks immediately after attention modules to extract hidden states. This enables the computation of a localized loss ${\mathcal{L}}_{\text{Block}}$ between the semi-quantized model and its auxiliary full precision counterpart, which is used to trigger a one-time localized backward pass per block. Here, to balance efficiency and accuracy in deriving importance scores, we define a block as an attention module as formulated in Equation 1, which encompasses multi-head self-attention along with Q/K/V/O projections and the residual stream. Additionally, we employ the block output mean square error (MSE) as the localized objective as
[…]
where $\mathbf{X}$} and $\hat{{\mathbf{X}}}$ are the full precision activations and activations with previous layers quantized, respectively. While it is theoretically possible to treat a full decoding layer as a block, doing so often incurs substantial memory overhead (Ding et al. 2023), rendering the method impractical on resource-constrained hardware.
Backward. The Backward module illustrates the internal mechanics within an attention layer. During this step, we perform a single pass of backpropagation on ${\mathcal{L}}_{\text{Block}}$ defined by Equation 8 and cache the gradients of each linear projection output (i.e., Q/K/V/O projections) to estimate token-level importance factors. We propose to use gradient-driven importance factors to quantify the contribution of each token, which is supported by Theorem 5.1.
Theorem 5.1.
The block-wise loss perturbation $\Delta{\mathcal{L}}_{\text{Block}}$ can be approximate by the first-order error as
$\Delta{\mathcal{L}}_{\text{Block}}\approx\Delta\theta{\mathbf{p}}^{(\Delta\theta),\top}+{\mathcal{O}}(|\theta|^{2})\approx\Delta{\mathbf{z}}{\mathbf{p}}^{(\Delta{\mathbf{z}}),\top},$ (9)
where $\theta\in{\mathbb{R}}^{D}$ and ${\mathbf{z}}\in{\mathbb{R}}^{Q}$ are the stacking weight being quantized and layer output, respectively.
The proof is provided in Appendix A. To this end, the gradient ${\mathbf{p}}^{(\Delta{\mathbf{z}})}$ measures the disturbance of layer output error on block loss, which can be an effective metric for token importance. Henceforth, we omit the superscript of ${\mathbf{p}}^{(\Delta{\mathbf{z}})}$ and reshape it to ${\mathbf{P}}\in{\mathbb{R}}^{C_{o}\times N}$ as our raw gradients.
Process Grad. Finally, the Process Grad module illustrates how we construct the importance factor matrix ${\mathbf{G}}\in{\mathbb{R}}^{N\times N}$ from the raw gradients $\mathbf{P}$}. Formally, the importance factor is defined as:
${\mathbf{G}}=\operatorname{Diag}\left(\left[\overline{\left|{\mathbf{P}}\right|}_{0},\,\overline{\left|{\mathbf{P}}\right|}_{1},\,\dots,\,\overline{\left|{\mathbf{P}}\right|}_{N-1}\right]\right),$ (10)
where the importance of $n$-th token is the $\ell_{1}$-norm of $n$-th column in gradients:
$\overline{\left|{\mathbf{P}}\right|}_{n}=\frac{1}{C_{o}}\sum_{i=0}^{C_{o}-1}\left|{\mathbf{P}}\right|_{i,n}.$ (11)
Note that importance factors are also computed for text tokens to preserve the relative contribution balance across modalities. These gradient-driven importance factors not only reflect inter-layer dependencies but also implicitly indicate how to adjust for cumulative errors introduced by previous quantized layers. Figure 4 visualizes importance factors of k_proj and v_proj output within a representative attention block. It is evident that the distribution of token salience exhibits distinct patterns across layers. For instance, several vision tokens are identified as salient in the k_proj output, whereas in the v_proj output, vision tokens are predominantly assigned low importance, indicating redundancy. This distribution highlights the diverse contribution of vision tokens to block-wise loss and further echoes the necessity of applying fine-grained, importance-aware calibration. To quantize the full model, we progressively alternate between local backpropagation and calibration, which better captures error propagation dynamics. The VLMQ is outlined in Appendix C.
6 Experiments
[…]
Qwen2.5-VL-7B-Instruct-INT3
Full Prec 83.20 94.72 58.55 52.98 84.40 88.26 70.62 83.05 76.97
GPTQ 62.44 89.33 42.31 27.33 77.70 79.39 66.58 78.38 65.43
GPTAQ 63.68 90.41 47.18 36.49 79.60 80.52 66.40 77.94 67.78
[…]
Qwen2-VL-2B-Instruct-INT3
Full Prec 72.61 89.35 40.27 43.59 76.60 74.23 61.79 79.38 67.23
GPTQ 58.84 74.55 33.25 30.83 61.50 65.13 50.37 73.49 56.00
GPTAQ 58.76 74.03 38.71 28.71 61.10 64.25 52.17 73.18 56.36
[…]
Qwen2-VL-7B-Instruct-INT3
Full Prec 81.44 93.89 57.30 56.09 80.90 85.55 69.26 82.02 75.81
GPTQ 72.80 88.35 47.76 27.90 72.40 78.26 63.02 78.28 66.10
GPTAQ 72.76 88.09 49.76 43.96 73.80 80.41 64.21 79.23 69.03
[…]
LLaVA-OneVision-7B-INT3
Full Prec 80.08 87.09 57.39 53.93 62.10 89.98 64.82 75.96 71.42
GPTQ 74.24 77.82 46.47 42.59 56.60 85.26 60.65 71.90 64.44
GPTAQ 75.72 78.07 46.48 42.94 58.30 85.97 61.84 72.69 65.25
[…]
Table 2: INT3 performance comparison across different quantization methods on eight benchmarks, which are ChartQA, DocVQA (Validation set), MME-RealWorld (English), MME-RealWorld (Chinese), OCRBench, ScienceQA, SeedBench 2 Plus, and TextVQA (Validation set), respectively. Best and second-best results are highlighted in bold and underline formats, respectively. indicates that we use the token-level $\ell_{2}$-norm of row gradients instead of $\ell_{1}$-norm as our importance factors.
[…]
Qwen2.5-VL-7B-Instruct-INT2g128
GPTQ 57.72 78.79 44.94 13.89 70.80 55.77 48.40 74.36 55.58
GPTAQ 53.48 74.82 39.38 3.60 67.70 2.95 17.83 73.79 41.69
[…]
Qwen2-VL-7B-Instruct-INT2g128
GPTQ 56.44 74.90 41.33 34.11 62.60 50.37 51.78 71.93 55.43
GPTAQ 56.08 72.57 37.91 33.80 61.30 59.70 51.25 67.98 55.07
[…]
LLaVA-OneVision-7B-INT2g128
GPTQ 61.08 62.62 40.29 31.35 48.60 67.88 51.25 67.29 53.80
GPTAQ 62.12 62.81 42.59 32.45 49.90 66.12 50.94 67.35 54.29
[…]
Table 3: INT2g128 performance comparison. indicates that we choose GPTQ as our precursor algorithm instead of GPTAQ.
Implementation Details
Models. We conduct experiments on open-source SOTA VLMs, including LLaVA-OneVision-0.5B/7B, Qwen2.5-VL-2B/7B/32B-Instruct, and Qwen2-VL-7B-Instruct. For the LLaVA-OneVision series, we select versions adopting Qwen2 as LM backbones and SigLIP-400M (Zhai et al. 2023) as their vision encoder.
Calibration. Following MBQ (Li et al. 2025a), we select the improved COCO Caption dataset introduced by ShareGPT4V (Chen et al. 2024b). We randomly sample 512 text-image pairs as our calibration set. Unless otherwise specified, we use the default dampening ratio $\lambda=0.01$ and enable act_order for all experiments. Other details can be found in Appendix D.
Evaluation. We undertake an extensive evaluation of quantized VLMs using multiple challenging vision-language tasks based on the LMMs-Eval framework (Zhang et al. 2024a). Specifically, we employ representative tasks including ChartQA (Masry et al. 2022), DocVQA (Mathew, Karatzas, and Jawahar 2021), MME-RealWorld (Zhang et al. 2024b), OCRBench (Liu et al. 2024), ScienceQA (Lu et al. 2022), SeedBench 2 Plus (Li et al. 2024a), and TextVQA (Singh et al. 2019), thereby comprehensively measuring the text recognition, visual perception, and visual reasoning capabilities of quantized models. Flash Attention (Dao et al. 2022) is enabled, and default configurations are used throughout all experiments.
Main Results
INT3 quantization. We assess the zero-shot performance of INT3 quantized models across eight vision-language benchmarks in Table 2. VLMQ demonstrates overall superiority compared to baseline approaches and achieves up to 1.01% average accuracy improvement. For Qwen2.5-VL-7B-Instruct-INT3, the proposed VLMQ outperforms its counterparts quantized by GPTQ and GPTAQ on all benchmarks except MME-RealWorld (English). Remarkably, it yields accuracy improvements of 1.64%, 1.40%, and 1.18% on ChartQA, MME-RealWorld (Chinese), and ScienceQA, respectively. For the LLaVA-OneVision-7B-INT3 model, we apply the token-level $\ell_{2}$-norm instead of the $\ell_{1}$-norm adopted in the other experiments, which we empirically find beneficial for performance enhancement.
INT2g128 quantization. We further evaluate the performance of INT2 group-wise quantized models with group_size=128 as shown in Table 3. While taking GPTAQ as our precursor algorithm generally yields satisfactory results under INT3 quantization, applying this configuration to INT2g128 leads to severe performance degradation. In particular, GPTAQ produces even worse results than GPTQ under this ultra-low-bit setting. Taking Qwen2-VL-7B-Instruct-INT2g128 as an example, GPTAQ achieves only 41.69% accuracy, significantly lower than the 55.58% achieved by GPTQ. Additionally, the Qwen2.5-VL-7B-Instruct-INT2g128 fails on MME-RealWorld (Chinese) and ScienceQA benchmarks, achieving only 3.60% and 2.95% accuracy, respectively. A plausible explanation is that, under extremely low-bit quantization, the residual error in GPTAQ propagated from previously quantized layers may be continuously accumulated and significantly amplified. This accumulated noise severely interferes with the dominant weight update term $\left(\hat{{\mathbf{w}}}_{q}-{\mathbf{w}}_{q}\right)/{\mathbf{H}}_{qq}^{-1}$, thereby compromising the undesirable quantization quality. To mitigate this issue, we choose GPTQ as our foundation algorithm. Evidently, we observe a notable performance improvement of VLMQ under the INT2g128 configuration. For instance, the MME-RealWorld (Chinese) benchmark exhibits a notable 16.45% accuracy gain for Qwen2.5-VL-7B-Instruct-INT2g128 compared with GPTQ.
Ablation Study
Bit Calib Data Bias Layer Acc
LLaVA-OneVision-7B
FP - - 71.42
INT3 WikiText None 65.01
COCO None 65.25
COCO Q/K/V Proj 64.73
COCO Q/K/V/O Proj 65.69
Qwen2-VL-7B-Instruct-INT3
FP - - 75.81
INT3 WikiText None 66.93
COCO None 69.03
COCO Q/K/V Proj 56.21
COCO Q/K/V/O Proj 69.42
Table 4: Ablation studies on LLaVA-OneVision-7B and Qwen2-VL-7B-Instruct.
We conduct step-by-step ablation studies on Qwen2-VL-7B-Instruction-INT3 and LLaVA-OneVision-INT3, as summarized in Table 4. It is evident that using WikiText alone as the calibration dataset leads to substantial performance degradation, underscoring the necessity of involving vision modality at quantization time. By incorporating token-level importance variance, VLMQ augments the Hessian matrix with importance-aware weighting, thereby improving the weight update step $\Delta{\mathbf{W}}$ and enhancing overall quantization quality. Notably, we find that applying our biased calibration strategy to o_proj is critical. We attribute this to the error accumulation caused by quantizing q_proj, k_proj, and v_proj. Leveraging enhanced Hessian information in o_proj allows effective error compensation within each residual block, ultimately improving quantization fidelity.
Quantization Efficiency
Model Size GPU Hours
GPTQ GPTAQ VLMQ
[…]
Table 5: Quantization time (in GPU hours) of 2B to 32B models using a single H100 80GB GPU. VLMQ incurs negligible overhead compared to baseline methods.
Table 5 presents the quantization time cost of VLMQ compared to baseline methods. As VLMQ is formally compatible with the GPTAQ formulation (Li et al. 2025b), it can seamlessly adopt efficiency-enhancing techniques such as Cholesky decomposition and residual decomposition. Because VLMQ incorporates only token-level importance, it results in minimal structural modifications to the optimization process. The primary additional overhead arises from a single local forward and backward pass per decoding layer, which introduces negligible latency in practice.
7 Conclusion and Limitation
We introduced VLMQ, a novel importance-aware PTQ framework for VLMs. We identified the vision over-representation that hinders the direct application of existing LLM PTQ methods to VLMs. Motivated by this insight, VLMQ explicitly leverages the enhanced Hessian yielded from the importance-aware objective. VLMQ effectively enhances the performance of quantized VLMs across various vision-language benchmarks, making large VLMs more practical for deployment under ultra-low-bit settings. In terms of limitations, our evaluation primarily focuses on image-text tasks. However, we believe VLMQ can generalize to broader applications, such as video understanding and language-only tasks, which we leave for future exploration.
Appendix
Appendix A Theoretical Proof
Proof of Equation 7
We begin by revisiting our importance-aware Lagrangian formulation, as presented in Equation 6, given by
[…]
To identify the local minima of the Lagrangian formulation, we differentiate with respect to $\Delta{\mathbf{w}}$ and the Lagrange multiplier $\lambda$, and solve the resulting equations by setting them to zero:
$\begin{cases}&\frac{\partial L}{\partial\Delta{\mathbf{w}}}=2\Delta{\mathbf{w}}\widetilde{{\mathbf{H}}}-2\widetilde{{\mathbf{r}}}\widetilde{{\mathbf{X}}}^{\top}+\lambda{\mathbf{e}}_{q}=0,\\ &\frac{\partial L}{\partial\lambda}=\Delta{\mathbf{w}}{\mathbf{e}}_{q}^{\top}+{\mathbf{w}}_{q}-\hat{{\mathbf{w}}}_{q}=0,\end{cases}$ (13)
which yields the solution for the first equation as
$\Delta{\mathbf{w}}\widetilde{{\mathbf{H}}}=-\lambda/2{\mathbf{e}}_{q}+\widetilde{{\mathbf{r}}}\widetilde{{\mathbf{X}}}^{\top}$. (14)
The following derivation closely follows the formal framework established in GPTAQ (Li et al. 2025b), with the key distinction being the substitution of the vanilla Hessian, residual, and activation matrix by their importance-aware counterparts, as defined in Equation 7.
Proof of Theorem 5.1
Theorem 5.1 establishes a connection between block-wise loss perturbation and layer output error. The proof begins with a Taylor expansion of the block-wise MSE loss with respect to the quantization noise. Specifically, the perturbation in the loss induced by quantized weights can be approximated as
$\Delta{\mathcal{L}}_{\text{Block}}$ $={\mathcal{L}}_{\text{Block}}\left(\theta+\Delta\theta\right)-{\mathcal{L}}_{\text{Block}}\left(\theta\right)$ (15)
$=\Delta\theta{\mathbf{p}}^{(\Delta\theta),\top}+{\mathcal{O}}(|\theta|^{2}),$ (16)
where $\theta\in{\mathbb{R}}^{D}$ is the stacking weight being quantized. By defining the layer output as ${\mathbf{z}}\in{\mathbb{R}}^{Q}$, we further derive the Equation 15 as
$\Delta{\mathcal{L}}_{\text{Block}}$ $\approx\Delta\theta{\mathbf{p}}^{(\Delta\theta),\top}$ (17)
$=\sum_{i=1}^{D}\Delta\theta_{i}\frac{\partial{\mathcal{L}}_{\text{Block}}}{\partial\theta_{i}}$ (18)
$=\sum_{i=1}^{D}\Delta\theta_{i}\left(\sum_{j=1}^{Q}\frac{\partial{\mathcal{L}}_{\text{Block}}}{\partial z_{j}}\frac{\partial z_{j}}{\partial\theta_{i}}\right)$ (19)
$=\sum_{j=1}^{Q}\frac{\partial{\mathcal{L}}_{\text{Block}}}{\partial z_{j}}\sum_{i=1}^{D}\left(\Delta\theta_{i}\frac{\partial z_{j}}{\partial\theta_{i}}\right)$ (20)
$=\sum_{j=1}^{Q}\Delta z_{j}\frac{\partial{\mathcal{L}}_{\text{Block}}}{\partial z_{j}}$ (21)
$=\Delta{\mathbf{z}}{\mathbf{p}}^{(\Delta{\mathbf{z}}),\top}.$ (22)
Proof done.
Appendix B Preliminary
Figure 5: VLM architecture.
Related Works
PTQ for LLMs. Post-training quantization (PTQ) (Lv et al. 2024; Chen et al. 2024a; Tian et al. 2024; Gong et al. 2024) has become a widely adopted technique for compressing (Li and Panda 2024; Wnag et al. 2024; Huang et al. 2025a) and accelerating large models (Hu et al. 2025; Huang et al. 2025c). For LLMs, methods such as AWQ (Lin et al. 2023) and SmoothQuant (Xiao et al. 2023) address activation outliers via smooth-based transformations. Hessian-guided approaches like OBS (Frantar and Alistarh 2022), GPTQ (Frantar et al. 2022), and GPTAQ (Li et al. 2025b) offer closed-form, layer-wise quantization solutions. More recently, rotation-based techniques (Ashkboos et al. 2024) enhance quantization-friendliness by reshaping parameter distributions, leading to improved low-bit performance. In parallel, reconstruction-based methods (Huang et al. 2024, 2025b) such as OmniQuant (Shao et al. 2023), TesseraQ (Li and Panda 2024), and CBQ (Ding et al. 2023) operate orthogonally to the above and can be integrated post-quantization for further refinement.
VLM Quantization Background
VLM architecture. As shown in Figure 5, the VLMs comprise three key components: a visual encoder, a vision-text projector, and a Language Model (LM) backbone. The visual encoders can be a pre-trained image encoder like a CLIP-style encoder (Radford et al. 2021), which is responsible for converting image/video input into vision tokens. Then they are fed into the MLP-based projector to align it with the LM embedding space. Together with encoded text tokens, they are further fed into the LM backbone. The LM backbone usually comprises multiple stacked decoding layers followed by an LM prediction head. The modality gap across hidden states in the word embedding space can be observed by PCA.
Given the multi-modal input ${\mathbf{X}}\in{\mathbb{R}}^{C_{i}\times N}$ of language model backbone, we can disassemble it based on different functionalities as
${\mathbf{X}}={\mathbf{X}}_{\texttt{sys}}\oplus{\mathbf{X}}_{\texttt{img}}\oplus{\mathbf{X}}_{\texttt{ins}}\oplus{\mathbf{X}}_{\texttt{ans}},$ (23)
where ${\mathbf{X}}_{\texttt{sys}}$, ${\mathbf{X}}_{\texttt{img}}$, ${\mathbf{X}}_{\texttt{ins}}$, and ${\mathbf{X}}_{\texttt{ans}}$ represent system prompt, vision, user instruction, and model response tokens, respectively. $\oplus$ signifies the concatenation operation. Without loss of generality, we can categorize tokens in different modalities and rewrite Equation 23 as
${\mathbf{X}}={\mathbf{X}}_{t}\oplus{\mathbf{X}}_{v},$ (24)
where text and vision token set ${\mathbf{X}}_{t}$ and ${\mathbf{X}}_{v}$ are
${\mathbf{X}}_{t}={\mathcal{S}}\{{\mathbf{X}}_{\texttt{sys}},{\mathbf{X}}_{\texttt{ins}},{\mathbf{X}}_{\texttt{ans}}\}$ (25)
${\mathbf{X}}_{v}={\mathcal{S}}\{{\mathbf{X}}_{\texttt{img}}\}$ (26)
Therefore, the Hessian information matrix can be written as
${\mathbf{H}}=\sum_{i=0}^{N_{t}-1}{\mathbf{x}}_{i}{\mathbf{x}}_{i}^{\top}+\sum_{j=0}^{N_{v}-1}{\mathbf{x}}_{j}{\mathbf{x}}_{j}^{\top}={\mathbf{H}}_{t}+{\mathbf{H}}_{v}$ (27)
where ${\mathbf{H}}_{t}$ and ${\mathbf{H}}_{v}$ represent the Hessian matrix formulated by text and vision tokens, respectively. Generally, the vision token count is much more than the text token count (i.e. $N_{v}\gg N_{t}$) and always exhibits high-degree of redundancy, leading to the overall feature implied by $\mathbf{H}$} biased towards vision modality.
Quantization basic. This paper focuses on uniform affine quantization. Quantization essentially maps high-precision (e.g., FP16/BF16) weights/activations to low-precision (e.g., INT8/INT4) formats to reduce memory footprint and achieve inference acceleration, which is given by
$\text{Quant: }{\mathbf{W}}_{\text{int}}=\operatorname{Clamp}\left(\Bigl{\lfloor}\frac{{\mathbf{W}}}{s}\Bigr{\rceil}+z,\,0,\,2^{B}-1\right),$ (28)
$\text{Dequant: }\hat{{\mathbf{W}}}=s\times\left({\mathbf{W}}_{\text{int}}-z\right),$ (29)
where ${\mathbf{W}}_{\text{int}}$ is in the INT-$B$ format and $\hat{{\mathbf{W}}}$ is the pseudo-quantized weight in full precision. The quantization parameters $s$ and $z$ are the quantization step size and zero-point, given by
$s$ $=\frac{\max\left({\mathbf{W}}\right)-\min\left({\mathbf{W}}\right)}{2^{B}-1}$ (30)
$z$ $=-\Bigl{\lfloor}\frac{\min\left({\mathbf{W}}\right)}{s}\Bigr{\rceil}.$ (31)
Hessian-based Quantization
OBQ calibrates the quantization of linear layers in models by iteratively 1) quantizing entries that introduce minimal loss, and 2) updating the remaining unquantized ones with the guidance of Hessian ${\mathbf{H}}={\mathbf{X}}{\mathbf{X}}^{\top}$. GPTQ (Frantar et al. 2022) extends this paradigm by formulating fixed-order entry quantization in a parallel manner, lazy block updating and Cholesky reformulation to achieve efficient calibration. GPTAQ (Li et al. 2025b) further identifies the critical error accumulation issue and proposes an asymmetric calibration principle to enhance the performance. Other Hessian-based variants (Kim et al. 2024; Edalati et al. 2025) turn to find a more accurate Hessian representation to preserve the quantized model capabilities.
Algorithm 1 VLMQ for one decoding layer
Input: Calibration input $\hat{{\mathbf{X}}}$, FP input $\mathbf{X}$}, and decoding layer ${\mathcal{W}}\leftarrow\{\texttt{Attn}(\cdot),\texttt{MLP}(\cdot)\}$.
[…]
 Cache ${\mathbf{G}}\leftarrow\texttt{Backward}({\mathcal{L}}_{\text{Block}})$ based on Equation 8.
for ${\mathbf{W}}_{\text{{name}}}\in{\mathcal{W}}$ do
  if $\texttt{name}\in\{\texttt{q\_proj,k\_proj,v\_proj,o\_proj}\}$ then
   Compute $\widetilde{{\mathbf{H}}}$, $\tilde{{\mathbf{r}}}$, and $\widetilde{{\mathbf{X}}}$ based on Equation 7.
[…]
  else
   Compute normal $\mathbf{H}$} and $\mathbf{r}$}.
   $\hat{{\mathbf{W}}}_{\texttt{name}}\leftarrow\text{GPTAQ}\left({\mathbf{W}}_{\texttt{name}},{\mathbf{H}},{\mathbf{r}},{\mathbf{X}}\right)$
  end if
end for
Appendix C VLMQ
The detailed algorithm is provided in Algorithm 1. For each decoding layer, we begin by caching the quantized input $\mathbf{X}$} and its full-precision counterpart $\hat{{\mathbf{X}}}$. Prior to quantizing the linear modules, we perform a block-wise forward and backward pass to obtain gradients of the Q/K/V/O projection outputs. The token-level importance factors are then computed from these gradients, as described in Equation 10. For quantizing the Q/K/V/O projections, we construct the importance-aware Hessian and apply the weight update rule defined in Equation 7. For the Up/Gate/Down projections, we formulate the vanilla Hessian and similarly apply the weight update rule in Equation 4.
Appendix D Experiments
Implementation Details
Calibration. We randomly sample 512 data samples as the calibration set. Each sample is structured according to Equation 23 and concatenated with the default prompt template. The input sequence length is adjusted based on the target model family: for Qwen2-VL and Qwen2.5-VL series, we truncate inputs to 512 tokens, while for LLaVA-OneVision series, the length is set to 986 tokens. To ensure semantic completeness, we discard samples whose vision token segments are truncated mid-way, as such instances result in incomplete symbolism. For quantization parameters, we adopt most settings from GPTAQ (Li et al. 2025b). Specifically, we enable act_order to enhance numerical stability. For group-wise quantization, we disable static_group to allow dynamic group-wise quantization parameter assignment.
Evaluation. We implement VLMQ using the Huggingface Transformers library (Wolf et al. 2019) on top of the PyTorch framework. For evaluation, we leverage the open-source LMMs-Eval toolkit (Zhang et al. 2024a), with minor adaptations tailored to our quantized settings. During evaluation, we observe that the default answer matching mechanism in LMMs-Eval is suboptimal for ScienceQA, especially under ultra-low-bit quantization scenarios. In particular, although the model often provides the correct answer, it may not strictly adhere to the required output format, namely, returning only the option letter. For instance, a response such as “The answer is B.” semantically matches the correct answer but fails to conform to the expected literal format (e.g., “B” or “B.”), leading to a false negative in accuracy computation. To mitigate this mismatch, we introduce a robust post-processing pipeline for automatic answer normalization and extraction. This process involves parsing the model’s textual response and identifying the most probable option letter based on regular expressions and contextual cues. The implementation details of this logic are provided in Listing 1. This enhancement ensures fairer evaluation and better reflects the true capability of quantized models on instruction-following tasks such as ScienceQA.
[…]
Qwen2.5-VL-32B-Instruct-INT3
GPTQ 43.88 88.26 51.77 47.76 76.50 87.22 64.87 73.88 74.20
GPTAQ 52.52 88.72 52.13 42.67 77.40 79.46 67.89 72.88 73.68
VLMQ 49.84 87.47 53.05 43.45 76.90 83.92 65.96 73.44 74.31
LLaVA-OneVision-0.5B-INT3
Full Prec 61.48 69.01 38.94 32.13 57.60 63.36 53.05 65.83 63.46
GPTQ 9.32 9.35 13.40 21.63 21.70 10.09 0.57 4.34 10.96
GPTAQ 2.88 7.70 9.20 11.27 20.70 5.35 1.23 9.22 9.17
VLMQ 8.04 16.40 20.42 15.08 29.20 7.90 3.03 12.77 14.86
Table 6: INT3 performance comparison on 0.5B/32B models.
Additional results. Table 6 presents additional results on 0.5B and 32B models. It is evident that VLMQ consistently outperforms baseline approaches. However, despite its effectiveness, the 0.5B model still exhibits a notable performance gap compared to its full-precision counterpart, highlighting the challenge of quantizing smaller VLMs under ultra-low-bit settings.
Listing 1 Additional string matching logic for ScienceQA
1# pred is the model response string
2if len(pred) >= 2 and pred[1] == ".":
3 result = 1.0 if pred[0] == target else 0.0
4 return {"exact_match": result}
5if ’A.’ in pred or ’A\"’ in pred or ’A in pred or pred[-1] == ’A’:
6 result = 1.0 if target == ’A’ else 0.0
7 return {"exact_match": result}
8if ’B.’ in pred or ’B\"’ in pred or ’B in pred or pred[-1] == ’B’:
9 result = 1.0 if target == ’B’ else 0.0
10 return {"exact_match": result}
11if ’C.’ in pred or ’C\"’ in pred or ’C in pred or pred[-1] == ’C’:
12 result = 1.0 if target == ’C’ else 0.0
13 return {"exact_match": result}
14if ’D.’ in pred or ’D\"’ in pred or ’D in pred or pred[-1] == ’D’:
15 result = 1.0 if target == ’D’ else 0.0
16 return {"exact_match": result}